# Leveraging the Bias-Variance Tradeoff in Quantum Chemistry for Accurate Negative Singlet-Triplet Gap Predictions: A Case for Double-Hybrid DFT

### v2

@@ 4.4 Application to heptazine derivatives
To probe the applicability of our bias-correction strategy, we focused on four DFIST candidates proposed in Ref. [4]. At the T1 geometry, the STGs of HzTFEX2, HzPipX2, HzTFEP2, and HzTFET2 at the EOM-CCSD, ADC(2), L-CC2, and SCS-ADC(2) levels exhibit smaller magnitudes[4] than those of the twelve triangular systems discussed above. This difference is attributed to the change in geometry, as STGs generally increase when moving away from the ground state minimum (S0) as pointed out from the context of excited-state nuclear dynamics[77] and the pseudo-Jahn–Teller effect [25]. Despite the larger molecular size of the heptazine derivatives compared to the triangular systems, local methods such as L-ADC(2) and L-CC2 provide significant speedups over their canonical counterparts, while dh-DFT calculations remain computationally feasible.
@@
Figure 6: Heptazine derivatives proposed as candidates for light emitters exhibiting delayed fluorescence from inverted singlet and triplet excited states (DFIST) in Ref. [4] are shown along with their computed STG values (in eV) from various methods. B2GP-PLYP and PBE-QIDH indicate that the results of the base methods B2GP-PLYP and PBE-QIDH are corrected with respect to their variants with $a_{x}=0.75$ and $a_{c}=0.55$; more details are available in the SI. Atoms in white||||black||||blue||||red||||cyan are H||||C||||N||||O||||F.
@@
For all four systems, geometry optimization was performed at the $\omega$B97X-D3/def2TZVP level (see SI for details), followed by single-point excited-state energy calculations. The STGs of these four systems, obtained from various methods, are presented in Figure 6. In the absence of a high-fidelity reference such as TBE for these systems, L-CC2 results are considered the most reliable. We used B2GP-PLYP (75,55) energies as references to scale the predictions from B2GP-PLYP (65,36) and similarly corrected PBE-QIDH (69,33) predictions using the (75,55) configuration. Notably, the application of the internal scaling does not require any additional high-level reference to identify the low-bias or low-variance domain in the parameter space of dh-DFAs. This enables a straightforward application of our internally referenced scaling procedure, which corrects systematic errors without requiring external high-level reference data.
@@
Among the results from various dh-DFAs, SCS-PBE-QIDH shows the best agreement with L-CC2, achieving an MAD of 0.014 eV. In contrast, bias-corrected results for B2GP-PLYP and PBE-QIDH are systematically more negative than the L-CC2 values, with MADs of 0.057 eV and 0.048 eV, respectively. While one might argue that simply shifting the results of these two DFAs could render the STGs negative, our correction is performed without requiring reference values from another method. As in the case of triangular benchmark systems (shown in Figure1), B2GP-PLYP exhibits a smaller bias compared to PBE-QIDH. Furthermore, for triangular systems, we observed that L-CC2’s error relative to the TBE is consistently shifted by --0.013 eV, whereas the errors of the (75,55) variant of the two dh-DFAs are centered around zero (albeit with a large spread), as shown in Figure 5. Consequently, compared to L-CC2, the bias-corrected dh-DFAs are expected to exhibit a non-vanishing MSE; indeed, the MSE values for B2GP-PLYP (75,55) and PBE-QIDH (75,55) are --0.057 eV and --0.048 eV, respectively. Compared to a more accurate reference such as CC3 (which is practically challenging for the four heptazine derivatives), our bias-corrected results may show better agreement than the other methods discussed. Moreover, it is possible to identify combinations of $a_{x}$ and $a_{c}$ that yield low bias and low variance relative to L-CC2, which can be further explored for extending the bias correction scheme to larger datasets lacking high-level reference data.
5 Conclusions
@@
First principles excited state modeling plays a crucial rule in the identification of new molecular light emitters with a negative S1-T1 energy gap. Historically, theoretical interest in such molecules existed for a long time in the context of Hund’s rule violation[78, 79, 80, 81, 82]. Yet, revived interest in them increased since their identification through first principles modeling[83, 84]. Although individual S1 and T1 energies are challenging to predict with high precision, the error cancellation inherent in STG calculations allows for improved accuracy, even when the reported gaps are less than 0.1 eV. Benchmarking twelve STG energy gaps in triangular molecules, we find that local wavefunction methods—specifically L-CC2 and L-ADC2—deliver mean errors of approximately 15 meV with correspondingly low standard deviations, providing both exceptional accuracy and the computational speed required for high-throughput screening of medium size molecules such as azaphenalenes with bulky substitutions.
@@
In parallel, we show that double-hybrid DFT approximations—though initially marked by high systematic errors—can be significantly improved through linear regression-based bias correction without depending on a high-level reference. For example, the PBE-QIDH method in its default configuration (69% exact exchange and 33% MP2-level correlation) exhibits an impressively low variance (9 meV) but high mean errors; by adjusting the parameters to 75% exchange and 55% correlation, the mean error is reduced to 4 meV, albeit with increased variance. We propose a bias correction strategy to leverage the bias-variance tradeoff intrinsic the dh-DFA formalisms. Using a suitable low-bias method as an internal reference, the approach corrects systematic errors in low-variance models, thereby enabling quantitatively reliable negative STG estimations. In the case of four heptazine derivatives, the SCS and SOS variants of both B2GP-PLYP and PBE-QIDH—though not subjected to bias correction—yield STG predictions in better agreement with L-CC2, suggesting that these variants naturally mitigate some of the bias in standard dh-DFT. However, bias correction remains valuable for further refining low-variance models, particularly when an appropriate low-bias reference is available.
@@
The purpose of this study is not to advocate different models for different properties but to leverage the intrinsic bias-variance tradeoff for accurate STG predictions. Open questions remain regarding the performance of these parameterized models for positive STGs and systems with charge-transfer excitations[85]. Robust application of our strategy requires reliable reference datasets and could benefit from further bias-variance analysis using alternative information-based metrics [86]. Overall, the findings of this study establish a mathematically sound framework for predictive modeling in first-principles molecular design, with promising implications for the development of novel light emitters and other applications.
6 Supplementary Information
@@
i) Dataset of S1, T1 and S1-T1 energies of twelve triangular molecules calculated with various methods and basis sets. ii) Figures S1–S3 present error metrics for cc-pVDZ, cc-pVTZ, and aug-cc-pVTZ basis sets; iii) Tables S1 presents optimal $a_{x}$ and $a_{c}$ along with the associated error metrics for different energies; iv) Tables S2–S3 present excited state energies of heptazine derivatives; v) Equilibrium coordinates of heptazine derivatives.
7 Data Availability
@@
8 Acknowledgments
We thank Prof. Denis Jacquemin for commenting on the ADC(2) and CC2 results of the benchmark triangular molecules. We acknowledge the support of the Department of Atomic Energy, Government of India, under Project Identification No. RTI 4007. All calculations have been performed using the Helios computer cluster, which is an integral part of the MolDis Big Data facility, TIFR Hyderabad (http://moldis.tifrh.res.in).
9 Author Declarations
9.1 Author contributions

### v3

@@ IV.5 Application to heptazine derivatives
To probe the applicability of our bias-correction strategy, we focused on four DFIST candidates proposed in Ref. Aizawa et al., 2022. At the T1 geometry, the STGs of HzTFEX2, HzPipX2, HzTFEP2, and HzTFET2 at the EOM-CCSD, ADC(2), L-CC2, and SCS-ADC(2) levels exhibit smaller magnitudesAizawa et al. (2022) than those of the twelve triangular systems discussed above. This difference is attributed to the change in geometry, as STGs generally increase when moving away from the ground state minimum (S0) as pointed out in the context of excited-state nuclear dynamicsKarak et al. (2024) and the pseudo-Jahn–Teller effect Majumdar et al. (2024). Despite the larger molecular size of the heptazine derivatives compared to the triangular systems, L-ADC(2) and L-CC2 provide significant speedups, enabling the calculation of their excited states.
@@
Figure 6: Heptazine derivatives proposed as candidates for light emitters exhibiting delayed fluorescence from inverted singlet and triplet excited states (DFIST) in Ref. Aizawa et al., 2022 are shown along with their computed STG values (in eV) from various methods. B2GP-PLYP and PBE-QIDH indicate that the results of the base methods B2GP-PLYP and PBE-QIDH are corrected with respect to their variants with $a_{x}=0.75$ and $a_{c}=0.55$; more details are available in the SI. Atoms in white||||black||||blue||||red||||cyan are H||||C||||N||||O||||F.
@@
For all four systems, we performed geometry optimization at the $\omega$B97X-D3/def2TZVP level (see SI for Cartesian coordinates), followed by single-point excited-state energy calculations. The STGs of these four systems, obtained from various methods, are presented in Figure 6. In the absence of a high-fidelity reference such as TBE for these systems, we consider L-CC2 results to be the most reliable. We used B2GP-PLYP (75,55) energies as an internal reference to scale the predictions from B2GP-PLYP (65,36) and similarly corrected PBE-QIDH (69,33) predictions using the (75,55) configuration. Notably, the application of the internal scaling does not require any additional high-level reference to identify the low-bias or low-variance domain in the parameter space of dh-DFAs. As these error metrics have been determined for the twelve benchmark systems. This enables a straightforward application of our internally referenced scaling procedure, which corrects systematic errors without requiring external high-level reference data.
@@
Among the results from various dh-DFAs, SCS-PBE-QIDH shows the best agreement with L-CC2, achieving an MAD of 0.014 eV. In contrast, the bias-corrected results for B2GP-PLYP and PBE-QIDH are systematically more negative than the L-CC2 values, with MADs of 0.057 eV and 0.048 eV, respectively. While one might argue that simply shifting the results of these two DFAs could render the STGs negative, our correction is performed without requiring reference values from the reference L-CC2 values. As in the case of triangular benchmark systems (shown in FIG 1), B2GP-PLYP exhibits a smaller bias compared to PBE-QIDH. Furthermore, for triangular systems, we observed that L-CC2’s error relative to the TBE is consistently shifted by --0.013 eV, whereas the errors of the (75,55) variant of the two dh-DFAs are centered around zero (albeit with a large spread), as shown in Figure 5. Consequently, when compared to L-CC2, the bias-corrected dh-DFAs are expected to exhibit a non-vanishing MSE; indeed, the MSE values for B2GP-PLYP (75,55) and PBE-QIDH (75,55) are --0.057 eV and --0.048 eV, respectively. Compared to a more accurate reference such as CC3 (which is practically challenging for the four heptazine derivatives), our bias-corrected results may show better agreement than the other methods discussed. Moreover, it is possible to identify combinations of $a_{x}$ and $a_{c}$ that yield low bias and low variance relative to L-CC2, which can be further explored for extending the bias correction scheme to larger datasets lacking high-level reference data such as TBE values.
V Conclusions
@@
First principles excited state modeling plays a crucial role in the identification of new molecular light emitters with a negative S1-T1 energy gap. Historically, theoretical interest in such molecules existed for a long time in the context of Hund’s rule violationToyota (1988); Borden, Iwamura, and Berson (1994); Gallup (1987); Koseki, Nakajima, and Toyota (1985); Toyota and Nakajima (1986). Yet, revived interest in them has increased since their identification through first-principles modeling, not so long ago de Silva (2019); Ehrmaier et al. (2019). Although individual S1 and T1 energies are challenging to predict with high precision, the error cancellation inherent in the definition of STG allows for improved accuracy, even when the magnitudes of the reported gaps are less than 0.1 eV. Benchmarking the STGs of twelve triangular molecules, we find that local wavefunction methods, specifically L-CC2 and L-ADC(2), deliver mean errors of approximately 15 meV with correspondingly low standard deviations, providing both exceptional accuracy and the computational speed required for high-throughput screening of medium-sized molecules such as azaphenalenes with bulky substitutions.
@@
In parallel, we demonstrate that double-hybrid DFT approximations, although initially characterized by high systematic errors, can be significantly improved through linear regression-based bias correction, without relying on a high-level reference. For example, the PBE-QIDH method in its default configuration (69% exact exchange and 33% MP2-level correlation) exhibits an impressively low variance (9 meV) but high mean errors. By adjusting the parameters to 75% exchange and 55% correlation, the mean error reduces to 4 meV, albeit with increased variance. We propose a bias correction strategy to leverage the bias-variance tradeoff inherent in the dh-DFA formalisms. Using a suitable low-bias method as an internal reference, the approach corrects systematic errors in low-variance models, thereby enabling accurate negative STG estimations. In the case of four heptazine derivatives, the SCS and SOS variants of both B2GP-PLYP and PBE-QIDH yield STG predictions in better agreement with L-CC2, suggesting that these variants naturally mitigate some of the bias in standard dh-DFT. However, bias correction remains valuable for further refining low-variance models, particularly when an appropriate low-bias reference is available.
@@
The purpose of this study is not to advocate for developing property-specific models, but to leverage the intrinsic bias-variance tradeoff in the dh-DFA formalism for accurate STG predictions. Open questions remain regarding the performance of these parameterized models for positive STGs and systems with charge-transfer excitations Casanova-Páez and Goerigk (2021b). Robust application of our strategy requires reliable reference datasets and could benefit from further bias-variance analysis using alternative information-based metrics Peverati (2021). Overall, the findings of this study establish a mathematically sound framework for predictive modeling in first-principles molecular design, with promising implications for the development of novel light emitters and other applications.
VI Supplementary Information
@@
i) Dataset of S1, T1 and S1-T1 energies of twelve triangular molecules calculated with various methods and basis sets. ii) Figures S1–S3 present error metrics for cc-pVDZ, cc-pVTZ, and aug-cc-pVTZ basis sets; iii) Tables S1 presents optimal $a_{x}$ and $a_{c}$ along with the associated error metrics for different energies; iv) Tables S2–S3 present excited state energies of heptazine derivatives; v) Equilibrium coordinates of heptazine derivatives. Excited state energies of the twelve benchmark systems determined with various methods and basis sets are provided as CSV files in the repository: https://github.com/moldis-group/triangulenes12.
VII Data Availability
@@
VIII Acknowledgments
We thank Prof. Denis Jacquemin for commenting on the ADC(2) and CC2 results of the benchmark triangular molecules, and Dr. Marcos Casanova-Páez for useful suggestions. We acknowledge the support of the Department of Atomic Energy, Government of India, under Project Identification No. RTI 4007. All calculations have been performed using the Helios computer cluster, which is an integral part of the MolDis Big Data facility, TIFR Hyderabad (http://moldis.tifrh.res.in).
IX Author Declarations
IX.1 Author contributions